# Error mitigation for logical circuits using decoder confidence

### v1

@@ C.2 Expectation Value Estimation
Appendix D Time overhead caused by aborting
Here, we calculate the time overhead required to obtain $M_{0}$ accepted shots based on the abort rate. The advantage of aborting is that it saves time by not letting a circuit-to-be-discarded complete. For an $N$-window circuit which aborts at some point throughout its execution, we want to know, on average, how far the execution proceeds before it is aborted. We define $p_{\text{abort}}(n)$ as the probability that the circuit aborts because of the $n^{\text{th}}$ window. This requires that the current window causes the abort, and additionally that none of the previous windows have caused it. Therefore, $p_{\text{abort}}(n)=\frac{\rho(1-\rho)^{n-1}}{f}$. The average window at which the abort happens (provided that it happens) is calculated as $\langle n\rangle=\sum_{i=1}^{N}np_{\text{abort}}(n)$. We divide this by $N$ to obtain the average percentage of circuit execution before abort: $\frac{1-(1-f)(1+N\rho)}{N\rho f}$.
Let $M_{0}$ be the desired number of shots for the quantum task, and $M$ the number of executed shots when the abort rate is $\rho$. They are related by $M=\frac{M_{0}}{1-f}$. Out of these $M$ executions, a fraction 1f1-f complete, thus contributing fully to the time duration, while the other $f$ fraction contribute an average of $\frac{\langle n\rangle}{N}$. Overall, the mean time overhead caused by aborting at rate $\rho$ becomes:
$\frac{1-f+f\frac{\langle n\rangle}{N}}{1-f}=\frac{f/N}{(1-f)[1-(1-f)^{\frac{1}{N}}]}.$ (11)
@@ Appendix E QCELS parameters
Based on numbers reported in this reference for the chosen algorithmic accuracy (the maximum number of Trotter steps involved in any quantum circuit and the average number of clocks per Trotter step), we deduced that the maximum number of decoding windows in any quantum circuit is $1.38\text{\times}{10}^{9}$. We use our own numerical predictions for the LER dependence on the code distance at a physical error rate $p_{\text{phys}}=10^{-3}$, rather than the reference $10^{-4}$. To ensure the noise is sufficiently suppressed, we identify the smallest code distance $d$ such that the mean LEP per circuit satisfies $\mu_{N,d}\leq 10^{-2}$, and that is $d=21$. To obtain a LEP distribution with the same mean as $(N_{\text{windows}},d)=($1.38\text{\times}{10}^{9}$,19)$, we used $(N_{\text{windows}},d)=($2.38\text{\times}{10}^{5}$,11)$.
We considered parameters $J=16$ and $N=5$ from Ref. [31], Sec. IV.B (note that $N$ quoted here is different from the usual meaning of $N$ as number of decoding windows throughout our work). Therefore, $N_{\text{values}}=2\times J\times N=160$ expectation values need to be estimated through different quantum circuits. Each circuit should be run $M_{0}$ times, which means that $N_{\text{values}}M_{0}=$16\,000$\$ quantum circuit runs occur in total.
@@ Appendix F Code distance influence on resources
To counteract the time overhead caused by aborting some circuit executions, there is an additional beneficial factor of $19/21$ because the duration of each logical operation scales at most linearly with $d$. Furthermore, the surface code qubit count is reduced to $(19/21)^{2}\approx 82\%$, leading to a total spacetime increase of $1.64\times(19/21)^{3}-1\approx 21\%$.
@@ Appendix G DCS-QEM performance metrics
In Fig. G.1, we present a different performance metric for the abort protocol and the MLE procedure to the one in the main text, namely, the absolute mean bias, called ’Bias’ for short in the y-axis labels of panels (b,e). This metric does not take variance into account, so we plot the latter in panels (c,f). In the limit of an infinite number of repetitions, the noiseless estimator bias would drop to exactly zero.
Figure G.1: Estimating $\langle Z\rangle_{\text{th}}=0.8$ on a noisy quantum computer, starting from a $7.7(2)\%$ LER in absence of error mitigation. (a,d) Empirical output distributions of four expectation value estimators, obtained from $N_{\text{rep}}=2\times 10^{5}$ repetitions. Dashed curves are normal distributions with the same mean and variance as the respective datasets. The noiseless estimator corresponds to an ideal quantum computer, whose output variation is solely caused by statistical Bernoulli fluctuations. The circuit discard fraction is $50\%$. (b,e) The performance of the four estimators plotted against resource overhead. For a given estimator and discard fraction, the bias is defined as $|\langle Z\rangle_{\text{th}}-\sum_{i=1}^{N_{\text{rep}}}\langle Z\rangle_{i}/N_{\text{rep}}|$. The overhead calculation is explained in Appendix D. (c,f) The variance of the estimator output distributions.

### v2

@@ C.2 Expectation Value Estimation
Appendix D Multiplicative Time Cost Caused by Aborting
Here, we calculate the multiplicative time cost required to obtain $M_{0}$ accepted shots based on the abort rate. The advantage of aborting is that it saves time by not letting a circuit-to-be-discarded complete. For an $N$-window circuit which aborts at some point throughout its execution, we want to know, on average, how far the execution proceeds before it is aborted. We define $p_{\text{abort}}(n)$ as the probability that the circuit aborts because of the $n^{\text{th}}$ window. This requires that the current window causes the abort, and additionally that none of the previous windows have caused it. Therefore, $p_{\text{abort}}(n)={\rho(1-\rho)^{n-1}}/{f}$. The average window at which the abort happens (provided that it happens) is calculated as $\langle n\rangle=\sum_{i=1}^{N}np_{\text{abort}}(n)$. We divide this by $N$ to obtain the average fraction of circuit executed before abort:
$\frac{\langle{n}\rangle}{N}=\frac{1-(1-f)(1+N\rho)}{N\rho f}.$ (12)
Let $M_{0}$ be the desired number of shots for the quantum task, and $M$ the number of executed shots when the abort rate is $\rho$. They are related by $(1-f)M=M_{0}$. Of these $M$ executions, a fraction 1f1-f complete, thus contributing fully to the time duration, while the other fraction $f$ contribute an average of $\frac{\langle n\rangle}{N}$. Overall, the mean multiplicative time cost caused by aborting at rate $\rho$ becomes:
$\omega:=\frac{1-f+f{\langle n\rangle}/{N}}{1-f}=\frac{f/N}{(1-f)[1-(1-f)^{{1}/{N}}]}.$ (13)
@@ Appendix E QCELS Parameters
Based on numbers reported in [31] for the chosen algorithmic accuracy (the maximum number of Trotter steps involved in any quantum circuit and the average number of clocks per Trotter step), we deduced that the maximum number of decoding windows in any quantum circuit is $1.38\text{\times}{10}^{9}$. We use our own numerical predictions for the LER dependence on the code distance at a physical error rate $p_{\text{phys}}=10^{-3}$, rather than the reference $10^{-4}$. To ensure the noise is sufficiently suppressed, we identify the smallest code distance $d$ such that the mean LEP per circuit satisfies $\mu_{N,d}\leq 10^{-2}$, and that is $d=21$. To obtain a LEP distribution with the same mean as $(N_{\text{windows}},d)=($1.38\text{\times}{10}^{9}$,19)$, we used $(N_{\text{windows}},d)=($2.38\text{\times}{10}^{5}$,11)$.
We considered parameters $J=16$ and $N=5$ from Ref. [31, Sec. IV.B] (note the meaning of $N$ here differs from the usual meaning of decoding window count used in this paper). Therefore, $N_{\text{values}}=2JN=160$ expectation values need to be estimated through different quantum circuits. Each circuit should be run $M_{0}$ times, which means that $N_{\text{values}}M_{0}=$16\,000$\$ quantum circuit runs occur in total.
@@ Appendix F Code Distance Influence on Resources
The surface code qubit count scales quadratically with distance $d$, and the duration of each logical operation scales at most linearly with $d$. Therefore the spacetime volume increase is $(d_{\text{start}}/d_{\text{targ}})^{3}\omega-1\approx 21\%$, where $\omega$ is the multiplicative time cost defined in Eq. 13.
@@ Appendix G DCS-QEM Performance Metrics
Figure G.1: (a,c) The same data as in Fig. 7b,d but analysed using a different performance metric: the bias, defined in Eq. 14. (b,d) The variance of the estimator output distributions.
In Fig. G.1, we present an alternative performance metric for the data in Fig. 7; namely, the absolute mean bias:
$\text{bias}:=\bigg|\langle Z\rangle_{\text{th}}-\frac{1}{N_{\text{rep}}}\sum_{i=1}^{N_{\text{rep}}}\langle Z\rangle_{i}\bigg|,$ (14)
which we call ‘Bias’ for short in the y-axis labels of panels (a,c). This metric does not account for variance, so we plot the latter in panels (b,d). In the limit of an infinite number of repetitions, the noiseless estimator bias drops to exactly zero.
Panels (a,c) show that MLE could achieve the target absolute mean bias ($0.633\times 10^{-2}$) for the application described in Section VI.1 without any time overhead.